OpenStudy (anonymous):

MEDALS!!!! What is the standard form of the equation? 3y – 7 = –5x A. 5x + 3y = 7 B. 3y – 7 = –5x C. 3y = –5x + 7 D. 5x + 3y – 7 = 0

OpenStudy (anonymous):

@Isaiah.Feynman

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

@one2345

OpenStudy (isaiah.feynman):

The standard form of a linear equation is \[Ax+By+C = 0\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

standard form is Ax + By = C 3y – 7 = –5x +7 +7 ____________________ 3y = -5x +7 +5x +5x ________________ 3y + 5x = 7

OpenStudy (anonymous):

Oh... I got the answer already.... but thanks.